help

What is the output of the code snippet below? int [ ] [ ] numarray = t 8, 7, 6, 1 0, 0, 0 li System.out.print (numarray [0] [0]) System.out.print (numary System.out.print (numarray to1 to array 1] to]); a) 00 b) 87 c) 80 d) 08
Answer #1

for the first program:

output is c) 80

in 2D array elements are arranged as follows:

numarray[0][0]=8, numarray[0][1]=7 numarray[0][2]=6

numarray[1][0]=0, numarray[1][1]=0, numarray[1][2]=0

for the second program

Ans is a) 2 4 7 11 16 22

because

in 1st iteration:

initially s=n=1. Inside loop s becomes equal to 2=s+n and n gets incremented by 1

and while condition get satisfied

in 2nd iteration: n=2 so s=2+2=4 and n gets incremented by 1 and while condition gets satisfied.

in 3rd iteration: n=3 so s=4+3=7 and n gets incremented by 1 and while condition gets satisfied.

in 4th iteration: n=4 so s=7+4=11 and n gets incremented by 1 and while condition gets satisfied.

in 5th iteration: n=5 so s=11+5=16 and n gets incremented by 1 and while condition gets satisfied.

in 6th iteration: n=6 so s=16+6=22 and n gets incremented by 1 and while condition gets dissatisfied

thus program terminates.